ttbone45:

What is the slope of the line that passes through (2, 12) and (4, 20)? (1 point) −4 2 4 −2

Angle:

Hi @ttbone45 Welcome to QuestionCove! for this problem, we can use the Slope formula \(slope=\huge\frac{y_2-y_1}{x_2-x_1}\) when we have two points \((x_1,y_1)\) and \((x_2,y_2)\)

Angle:

if \(x_1=2\), \(y_1=12\), \(x_2=4\), and \(y_2=20\) you can plug it into the formula and simplify

Angle:

@ttbone45 if you are still stuck or confused, just let me know

Angle:

\(\huge \frac{20-12}{4-2}\) = ?

koolkidjay:

https://www.desmos.com/calculator use this it will tell u hope this helps